OpenStudy (anonymous):

Which is a factor of 6x2 + 33x – 63 ?

OpenStudy (anonymous):

would it be 3?

OpenStudy (anonymous):

sorry to bother everyone but i need help and fast!! can somebody please help mer!? study guide.. Solve for x: 2800=400*2¹⁰ˣ Check all that apply. A. x=log(7)/10 B. x=log(7)/10*ln(2) C. x=ln(7)/10*ln(2) D. x=ln(7)/10

OpenStudy (anonymous):

^^^^^??????

OpenStudy (anonymous):

$$\ \ \ \ 6x^2+33x-63\\=3(2x^2+11x-21)\\=3(2x^2+14x-3x-21)\\=3(2x(x+7)-3(x+7))\\=3(2x-3)(x+7)$$

hero (hero):

I think @oldrin.bataku has posted enough hints

OpenStudy (anonymous):

Okay... now which of those are factors we've found?

hero (hero):

Let @FlyinSolo_424 figure it out

hero (hero):

You're guessing

hero (hero):

You're right, 2x - 3 is not an answer choice. I think you should make an attempt at figuring out which answer choice it is.

hero (hero):

There's a difference between helping someone and being a jerk.

hero (hero):

Mind explaining why 3 isn't a possibility?

OpenStudy (anonymous):

\(3\) is the answer. The factors are what are being multiplied to yield the polynomial; in this case, we have: \(3, 2x-3,x+7\).

hero (hero):

Oh I see, @FlyingSolo_424, you want someone to just GIVE you the answer. I get it now.

hero (hero):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Thank you @oldrin.bataku

hero (hero):

@FlyinSolo_424, why did you delete your comments?

OpenStudy (anonymous):

cool

hero (hero):

I'm saddened that you believe I wasn't trying to help you. I was trying to help you figure it out on your own by leading you to the correct result. Sorry if you thought otherwise. And also, it wasn't necessary to block me.

OpenStudy (anonymous):

Well thats how it came off. Thays what people usually do lead me through it. One in specific person some bicycle guy. and I felt it was because you wouldnt leave me alone

hero (hero):

I'm sorry, but you lost me. I come to help you, and then you want me to leave you alone? How is that being nice?

OpenStudy (sirm3d):

sometimes it is appropriate to let the person helping her do his part. we come in only when needed.

hero (hero):

Yes, and I intervened because I thought that @oldrin.bataku was going to give you the answer and I just wanted to make sure you got a chance to figure it out yourself first.

OpenStudy (sirm3d):

just sit and wait. watch. read. relax. we're not the only ones who can do math right.

OpenStudy (anonymous):

Not everyone understands what is really being asked until you explain a trial problem out. I'm not a fan of spoon-feeding but learning by example isn't terrible... unless they don't actually learn.

hero (hero):

I know that @oldrin.bataku. That's why I didn't say anything to you initially.

OpenStudy (anonymous):

Because You kinda came off as rude. As you could tell i still didnt know the anwser when he wrote it out and you didnt see me write... "so whats the answer?" I was asking questions and he was helping and you jumped to conclusions.

hero (hero):

I wasn't trying to be rude. I was trying to help. I think maybe you are impatient and get upset really easily. Not everything is as it seems. I stated the reason why I intervened earlier. Let's quell this now and move on please.

OpenStudy (anonymous):

Take your own advice . How are you to judge me over a question? No I just hate when people assume . Making assumptions is probley the quickest way to make me mad.. especially when no knowing the situation. I think whats best in both of our favor is to leave eachother alone. So Please... Leave me alone.